# Probing entanglement dynamics and topological transitions on noisy intermediate-scale quantum computers

### v1

@@ I Introduction
Several studies have demonstrated the potential of digital quantum computing for quantum many-body simulations [12]. In particular, various entanglement properties, such as entanglement spectra [13], symmetry-resolved entropies [14], the second-order Rényi entropy [15, 16], for both many-body ground states and dynamical states have been actively explored on current quantum computers. There is also an increasing interest in probing topological properties with NISQ devices; recent works include measuring string order parameters crossing symmetry protected topological phase transitions [17], simulating chiral topological dynamics [18], and probing fractional quantum Hall states [19].
@@ II The Model and quench protocols
where $m$ indexes $N$ unit cells that are pairs of two sites, one on sublattice $A$ and one on $B$. The operators $a^{\dagger}_{m}$($b^{\dagger}_{m}$) creates a spinless fermion at the $A$($B$)-sublattice site in the $m$th unit cell, and $\mathcal{N}$ is the the number of fermions. The constants $2J$ and $2J^{\prime}$ are regarded as the intracell and intercell hopping amplitudes. The boundary term is affected by the parity of the number of fermions [26]; depending on evenness or oddness of $\mathcal{N}$, the boundary conditions of the fermionic chain are antiperiodic or periodic.
@@
with $\phi(k)=\tan^{-1}(d_{k}^{y}/d_{k}^{x})$. The dispersion relation in Eq.( 6) describes a band structure with a gap $\Delta_{E}=4\left|{J-J^{\prime}}\right|$ at $k=\pm\pi$, as long as $J\neq J^{\prime}$. In the fully dimerized cases ($J>0$, $J^{\prime}=0$, or vice versa $J=0$, $J^{\prime}>0$), the band structure becomes flat, i.e. independent of the wavenumber $k$; in this flat-band limit, there is also no difference between the periodic and antiperiodic boundary conditions. The topological properties of the system are encoded in the closed curve that the vector $\mathbf{d}_{k}$ traces out as $k$ changes from 00 to $2\pi$. For $J<J^{\prime}$, the curve encloses the origin $\mathbf{d}_{k}=\mathbf{0}$ and the system is in the topological phase; for $J>J^{\prime}$, the origin is completely outside the closed curve, corresponding to a trivial phase. At the critical point $J=J^{\prime}$ where the band gap closes, the curve passes through the origin. The number of times that the closed curve goes around the origin is referred to as the winding number $\nu$, defining a topological invariant. Other useful tools to characterize topological properties include entanglement measures [28, 21, 22] and the Berry phase [29, 30, 31, 32].
@@
in terms of the standard basis of $z$ spin components: $\sigma^{z}\left|{0}\right\rangle=\left|{0}\right\rangle$, $\sigma^{z}\left|{1}\right\rangle=-\left|{1}\right\rangle$, and (ii) a fully dimerized stated described by a product of intracell singlets written as
$\left|{\psi_{\text{S}}}\right\rangle=\bigotimes_{j=1}^{N}\frac{1}{\sqrt{2}}\left(\left|{0_{2j-1}1_{2j}}\right\rangle-\left|{1_{2j-1}0_{2j}}\right\rangle\right)\,.$ (9)
@@ II.2 Twist operator
in terms of the ground state of the XXZ spin chain can be seen as an indicator of a gapped or gapless excitation spectrum, depending on whether $z^{(q)}_{L}$ is nonvanishing or zero in the thermodynamic limit $L\to\infty$. Furthermore, the twist order parameter was generalized to higher spin cases and is capable of detecting the patterns of valence-bond-solid-like dimerization in the state $\left|{\psi}\right\rangle$ [27], even for random and inhomogeneous systems [38]. The real part of $z_{L}$ is negative/positive for a dimerized state with strong links connecting adjacent sites $(i,i+1)$ with even/odd $i$ (i.e. intercell/intracell links).
Here we apply the twist operator to the dynamical states in our quench protocols. As shown in Fig. 1(b) for $L=16$, the twist order parameter oscillates in the region $\text{Re}[z_{L}]\leq 0$ for a quench from the Néel state $\left|{\psi_{\text{N}}}\right\rangle$, while it oscillates in the region $\text{Re}[z_{L}]>0$ for a quench from the initial state $\left|{\psi_{\text{S}}}\right\rangle$. The oscillation of $\text{Re}[z_{L}]$ for the case of $\left|{\psi_{\text{N}}}\right\rangle$ indicates that the strongest dimerization occurs at $t^{*}_{\text{N}}$ when the maximum entanglement forming between nearest neighboring cells; for the case of $\left|{\psi_{\text{S}}}\right\rangle$, the system starts from a fully dimerized state with intracell links and returns to this strongest dimerization after a period of $2t^{*}_{\text{S}}$.
@@ III.2 The Rényi entanglement entropy
Figure 6: Results for the half-chain entanglement entropy for a chain with $L=8$ and PBC after quenches from the initial states $\left|{\psi_{\text{N}}}\right\rangle$ (a) and $\left|{\psi_{\text{S}}}\right\rangle$ ((b) and (c)), simulated on the IBM device ibmq_montreal. The raw data in (a) and (b) were obtained using $N_{U}=100$ sets of random unitaries, while for (c) $N_{U}=200$ was used. The data set in (a) with label ’shifted’ was obtained by shifting the raw data set by a constant to align the entropy at $t=0$ with the exact value $S(0)=0$. The results for $\left|{\psi_{\text{S}}}\right\rangle$ in (c) after applying error mitigation using the global depolarizing ansatz show an excellent match with the exact solution.
@@
where $I$ denotes a subsystem of $N_{I}$ qubits. The error rate $p_{\text{tot}}$ can be extracted from $\text{Tr}[\rho^{2}]$ when the full density matrix for a pure state is considered since $\text{Tr}[\rho^{2}_{\text{exact}}]=1$. We have determined $p_{\text{tot}}$ for each time step to carry out the error mitigation for the time-dependent data. In addition, the mitigated data in Fig. 6 are adjusted to align the minimum to zero to correct negative entropies in the minima. The results after error mitigation have clearly achieved quantitative agreement with the exact solutions except that the largest amplitudes of the oscillations in Fig. 6 (a) and (b) are lower than the theoretical value.
@@
The deviation from the exact solutions can have more than one source. For example, the accuracy of the estimated purity by randomized measurements depends crucially on the the number of random unitaries and the number of measurements per random unitary [43], as implied in Eq. (21). The results in Fig. 6 with $N_{U}=200$ after applying error mitigation have clearly achieved an excellent match with the exact data, while the mitigated results for the cases using $N_{U}=100$ are merely passable. Another major source of error for our case stems from the PBC imposed on the model, which requires connectivity between the first and last qubits via ancilla qubits and additional gates, thereby leading to higher error rates. For example, the circuit depth for the simulations of the chains with PBC is 40 (45) for the initial state $\left|{\psi_{\text{N}}}\right\rangle$ ($\left|{\psi_{\text{S}}}\right\rangle$), while the depth is reduced to 15 (19) when open chains with OBC are considered. We can indeed reproduce the results for $L=8$ with higher accuracy by using a segment of bulk qubits (from the 3rd to the 6th qubit) in an open chain as the subsystem considered for the EE. The results for the open chain displayed in Fig. 7 largely agree quantitatively with the exact data, even without applying error mitigation. After applying the global depolarizing error mitigation, the overall shape of the EE as a function of time, in particular around the entropy valleys, is corrected. However, we need to apply a constant shift to align the entropy at $t=0$ (the minimum entropy) to zero.
@@ III.3 The twist operator
The twist operator defined in Eq. (14) is equivalent to a product of $\bm{R_{z}}$ rotation operators: $\otimes_{j=1}^{L}\bm{R_{z}}(\theta_{j})$ with $\theta_{j}=-i4\pi j/L$ for the angle of rotation on the $j$th-qubit. In practice, we can circumvent the accumulated errors resulting from the series of rotation operators by directly measuring $\otimes_{j=1}^{L}\bm{Z}$ for each instant of time, a straightforward process in the computational basis. Given a set of measured bitstrings $\{\mathbf{s}:\mathbf{s}=s_{1}s_{2}\cdots s_{L}\text{ with }s_{j}=0$ or 1}1\}1 } with occurring probabilities $P(\mathbf{s})$, the twist order parameter $z_{L}$ (defined in Eq. (15)) is then obtained via postprocessing using
@@
The unmitigated data for $z_{L}$ in Fig. 10, both for the initial states $\left|{\psi_{\text{N}}}\right\rangle$ and $\left|{\psi_{\text{S}}}\right\rangle$, display oscillations with periods identical to those of the exact results, but have considerably smaller amplitudes. We have applied a simple postselection method by only considering bitstrings with an equal number of ’0’s and ’1’s, in accordance with particle number (magnetization) conservation requirements. We find an increase in amplitudes for the data after postselecting; in particular, the results for quenches starting from the Néel state $\left|{\psi_{\text{N}}}\right\rangle$ have shown good quantitative agreement with exact diagonalization results.
@@
for a set of bitstrings $\{\mathbf{s}\}$ with occurring probabilities $P(\mathbf{s})$. The Berry phase is then the argument of this complex amplitude: $\gamma_{L}=\text{Arg}[z_{N}^{(2)}]$. We only consider the case for the initial state $\left|{\psi_{\text{N}}}\right\rangle$ since the case for $\left|{\psi_{\text{S}}}\right\rangle$ has zero Berry phase under the quench dynamics. We have used the same bitstrings from the IBM device for $z_{L}$ and the same postselected bitstrings to calculate the amplitude $z^{(2)}_{N}$. In comparison to the twist order parameter $\text{Re}[z_{L}]$, the Berry phase, as the argument of $z^{(2)}_{N}$, appears to be more robust against noise. As shown in Fig. 11, the Berry phase obtained from the IBM Q device after error mitigation is in perfect agreement with exact diagonalization results.
@@ IV Summary and discussion
The LSM twist operator applied in our simulations can be used to formulate many quantities and concepts, such as order parameters to characterize gapped phases [27], Berry phases [37], polarization [37] and localization length [39], that are of fundamental importance in quantum many-body physics. Recently, the applications of the LSM twist operator have also been generalized to open quantum systems [48, 49]. The twist operator is diagonal in the computaional basis on quantum computers. This property allows us to calculate its related quantities on classical computers by processing the bit-strings and their corresponding probabilities from quantum device measurements, without incurring additional overhead in quantum hardware. The method makes many interesting applications of the twist operator amenable to simulations on NISQ devices.

### v2

@@ I Introduction
Several studies have demonstrated the potential of digital quantum computing for quantum many-body simulations [12]. In particular, various entanglement properties, such as entanglement spectra [13], symmetry-resolved entropies [14], the second-order Rényi entropy [15, 16], for many-body ground states and dynamical states have been actively explored on current quantum computers. There is also an increasing interest in probing topological properties with NISQ devices; recent works include measuring string order parameters crossing symmetry protected topological phase transitions [17], simulating chiral topological dynamics [18], and probing fractional quantum Hall states [19].
@@ II The Model and quench protocols
where $m$ indexes $N$ unit cells that are pairs of two sites, one on sublattice $A$ and one on $B$. The operators $a^{\dagger}_{m}$($b^{\dagger}_{m}$) creates a spinless fermion at the $A$($B$)-sublattice site in the $m$th unit cell, and $\mathcal{N}$ is the the number of fermions. The constants $2J$ and $2J^{\prime}$ are regarded as the intracell and intercell hopping amplitudes, respectively. The boundary term is affected by the parity of the number of fermions [26]; depending on evenness or oddness of $\mathcal{N}$, the boundary conditions of the fermionic chain are antiperiodic or periodic.
@@
with $\phi(k)=\tan^{-1}(d_{k}^{y}/d_{k}^{x})$. The dispersion relation in Eq.( 6) describes a band structure with a gap $\Delta_{E}=4\left|{J-J^{\prime}}\right|$ at $k=\pm\pi$, as long as $J\neq J^{\prime}$. In the fully dimerized cases ($J>0$, $J^{\prime}=0$, or vice versa $J=0$, $J^{\prime}>0$), the band structure becomes flat, i.e. independent of the wavenumber $k$. In this flat-band limit, there is also no difference between the periodic and antiperiodic boundary conditions. The topological properties of the system are encoded in the closed curve that the vector $\mathbf{d}_{k}$ traces out as $k$ changes from 00 to $2\pi$. For $J<J^{\prime}$, the curve encloses the origin $\mathbf{d}_{k}=\mathbf{0}$ and the system is in the topological phase; for $J>J^{\prime}$, the origin is completely outside the closed curve, corresponding to a trivial phase. At the critical point $J=J^{\prime}$ where the band gap closes, the curve passes through the origin. The number of times that the closed curve goes around the origin is referred to as the winding number $\nu$, defining a topological invariant. Other useful tools to characterize topological properties include entanglement measures [28, 21, 22] and the Berry phase [29, 30, 31, 32].
@@
in terms of the standard basis of $z$ spin components: $\sigma^{z}\left|{0}\right\rangle=\left|{0}\right\rangle$, $\sigma^{z}\left|{1}\right\rangle=-\left|{1}\right\rangle$, and (ii) a fully dimerized state described by a product of intracell singlets written as
$\left|{\psi_{\text{S}}}\right\rangle=\bigotimes_{j=1}^{N}\frac{1}{\sqrt{2}}\left(\left|{0_{2j-1}1_{2j}}\right\rangle-\left|{1_{2j-1}0_{2j}}\right\rangle\right)\,,$ (9)
@@ II.2 Twist operator
in terms of the ground state of the XXZ spin chain can be seen as an indicator of a gapped or gapless excitation spectrum, depending on whether $z^{(q)}_{L}$ is nonvanishing or zero in the thermodynamic limit $L\to\infty$. Furthermore, the twist order parameter was generalized to higher spin cases and is capable of detecting the patterns of valence-bond-solid-like dimerization in the state $\left|{\psi}\right\rangle$ [27], even for random and inhomogeneous systems [35]. The real part of $z_{L}$ is negative/positive for a dimerized state with strong links connecting adjacent sites $(j,j+1)$ with even/odd $j$ (i.e. intercell/intracell links).
Here we apply the twist operator to the dynamical states in our quench protocols. As shown in Fig. 1(b) for $L=16$, the twist order parameter oscillates in the region $\text{Re}[z_{L}]\leq 0$ for a quench from the Néel state $\left|{\psi_{\text{N}}}\right\rangle$, while it oscillates in the region $\text{Re}[z_{L}]>0$ for a quench from the initial state $\left|{\psi_{\text{S}}}\right\rangle$. The oscillation of $\text{Re}[z_{L}]$ for the case of $\left|{\psi_{\text{N}}}\right\rangle$ indicates that the strongest dimerization occurs at $t^{*}_{\text{N}}$ when the maximum entanglement forming between nearest neighboring cells; for the case of $\left|{\psi_{\text{S}}}\right\rangle$, the system starts from a fully dimerized state with intracell links and returns to this strongest dimerization after a period of $\pi/2$.
@@ III.2 The Rényi entanglement entropy
Figure 6: Results for the half-chain entanglement entropy for a chain with $L=8$ and PBC after quenches from the initial states $\left|{\psi_{\text{N}}}\right\rangle$ (a) and $\left|{\psi_{\text{S}}}\right\rangle$ (b), simulated on the IBM device ibm_montreal. The results for $\left|{\psi_{\text{S}}}\right\rangle$ after applying error mitigation using the global depolarizing ansatz show an excellent match with the exact solution.
@@
where $I$ denotes a subsystem of $N_{I}$ qubits. The error rate $p_{\text{tot}}$ can be extracted from $\text{Tr}[\rho^{2}]$ when the full density matrix for a pure state is considered since $\text{Tr}[\rho^{2}_{\text{exact}}]=1$. We have determined $p_{\text{tot}}$ for each time step to carry out the error mitigation for the time-dependent data. In addition, the mitigated data in Fig. 6 are adjusted to align the minimum to zero to correct negative entropies in the minima. The results for $\left|{\psi_{\text{S}}}\right\rangle$ after applying error mitigation clearly show an excellent match with the exact data, while the mitigated results for $\left|{\psi_{\text{N}}}\right\rangle$ with global depolarizing ansatz are merely passable.
@@
The deviation from the exact solutions can have more than one source. For example, the accuracy of the estimated purity by randomized measurements depends crucially on the the number of random unitaries and the number of measurements per random unitary [41], as implied in Eq. (21). However, the primary source of error for our case may stem from the PBC imposed on the model, which requires connectivity between the first and last qubits via ancilla qubits and additional gates, thereby leading to higher error rates. For example, the circuit depth for the simulations of the chains with PBC is 40 (45) for the initial state $\left|{\psi_{\text{N}}}\right\rangle$ ($\left|{\psi_{\text{S}}}\right\rangle$), while the depth is reduced to 15 (19) when open chains with OBC are considered. We can indeed reproduce the results for $L=8$ with higher accuracy by using a segment of bulk qubits (from the 3rd to the 6th qubit) in an open chain as the subsystem considered for the EE. The results for the open chain displayed in Fig. 7 largely agree quantitatively with the exact data, even without applying error mitigation. After applying the global depolarizing error mitigation, the overall shape of the EE as a function of time, in particular around the entropy valleys, is corrected. However, we need to apply a constant shift to align the entropy at $t=0$ (the minimum entropy) to zero.
@@ III.3 The twist operator
The twist operator defined in Eq. (14) is equivalent to a product of $\bm{R_{z}}$ rotation operators: $\otimes_{j=1}^{L}\bm{R_{z}}(\theta_{j})$ with $\theta_{j}=-2\pi j/L$ for the angle of rotation on the $j$th-qubit. In practice, we can circumvent the accumulated errors resulting from the series of rotation operators by directly measuring $\otimes_{j=1}^{L}\bm{Z}$ for each instant of time, a straightforward process in the computational basis. Given a set of measured bitstrings $\{\mathbf{s}:\mathbf{s}=s_{1}s_{2}\cdots s_{L}\text{ with }s_{j}=0$ or 1}1\}1 } with occurring probabilities $P(\mathbf{s})$, the twist order parameter $z_{L}$ (defined in Eq. (15)) is then obtained via postprocessing using
@@
The unmitigated data for $z_{L}$ in Fig. 10, both for the initial states $\left|{\psi_{\text{N}}}\right\rangle$ and $\left|{\psi_{\text{S}}}\right\rangle$, display oscillations with periods identical to those of the exact diagonalization (ED) results, but have considerably smaller amplitudes. We have applied a simple postselection method by only considering bitstrings with an equal number of ’0’s and ’1’s, in accordance with particle number (magnetization) conservation requirements. We find an increase in amplitudes for the data after postselecting; in particular, the results for quenches starting from the Néel state $\left|{\psi_{\text{N}}}\right\rangle$ have shown good quantitative agreement with ED results.
@@
for a set of bitstrings $\{\mathbf{s}\}$ with occurring probabilities $P(\mathbf{s})$. The Berry phase is then the argument of this complex amplitude: $\gamma_{L}=\text{Arg}[z_{N}^{(2)}]$. Here we only consider the case for the initial state $\left|{\psi_{\text{N}}}\right\rangle$ since the case for $\left|{\psi_{\text{S}}}\right\rangle$ has zero Berry phase in the quench dynamics. We have used the same bitstrings from the IBM device for $z_{L}$ and the same postselected bitstrings to calculate the amplitude $z^{(2)}_{N}$. In comparison to the twist order parameter $\text{Re}[z_{L}]$, the Berry phase, as the argument of $z^{(2)}_{N}$, appears to be more robust against noise. As shown in Fig. 11, the Berry phase obtained from the IBM Q device after error mitigation is in perfect agreement with exact diagonalization results.
@@ IV Summary and discussion
The LSM twist operator applied in our simulations can formulate various quantities and concepts fundamental to quantum many-body physics, such as order parameters to characterize gapped phases [27, 46], Berry phases [36], polarization [36] and localization length [37]. Recently, the applications of the LSM twist operator have also been generalized to open quantum systems [47, 48]. The twist operator is diagonal in the computational basis on quantum computers. This property allows us to calculate its related quantities on classical computers by processing the bit-strings and their corresponding probabilities from quantum device measurements, without incurring additional overhead in quantum hardware. The method makes many interesting applications of the twist operator amenable to simulations on NISQ devices.